OpenStudy (anonymous):

The number of tickets sold each day for an upcoming performance of Handel`s messiah is given by N(x)=-0.4x^2+8x+11, where x is the number of days since the concert was first announced. When will daily ticket sales peak, and how many tickets will be sold that day? I know the sales will peak at day 10. What formula do I use to figure out how many tickets will be sold?

OpenStudy (ybarrap):

N(10)

OpenStudy (anonymous):

so N(10)=-0.4x^2+8x+11?

OpenStudy (anonymous):

x=-b/(2a)=-8/(2*-.4)=10 so day 10

OpenStudy (anonymous):

N(10=-0.4(10)^2+8(10)+11=51 tickets

OpenStudy (anonymous):

remember this is just the graph of a parabola opening downward, the max will occur at the x-value of the vertex, and the value of the max will be the y-value which represents number of tickets in this problem

OpenStudy (anonymous):

@cali once you get \[-\frac{b}{2a}=10\] you replace x by 10 everywhere you see it in the expression to get \[-.4\times 10^2+8\times 10+11=-.4\times 100+80+11=-40+80+11=51\]

OpenStudy (anonymous):